OpenStudy (anonymous):

When invested at an annual interest rate of 8%, an account earned $336 of simple interest in one year. How much money was originally invested in the account?

OpenStudy (anonymous):

ok so we have r = 8 %, t = 1 year and I = 336 and we need to find P

OpenStudy (anonymous):

we have the formula I = PRT/100 that is 336 = .08 P

OpenStudy (anonymous):

divide by .08 on both sides, we have p = 336 /.08 that is 4200

OpenStudy (anonymous):

0.08x = 336 x = 336/0.08 x= 4200$

OpenStudy (anonymous):

thank you!! @sandyjayaram @zscdragon

OpenStudy (anonymous):

np :)